OpenStudy (anonymous):

If 9:x = x:4, then x is 6. True False

geerky42 (geerky42):

is 9:6 equal to 6:4?

OpenStudy (anonymous):

the problem can be rewritten as: 9/x = x/4 cross multiply and u get: x^2 = 36 solve for x if u cant solve for x from that point

OpenStudy (anonymous):

can* sorry

OpenStudy (anonymous):

starting from where Reaper left off... x^2 = 36 x = sq rt 36 x = 6 True

OpenStudy (anonymous):

the answer is true:) thanks for the help:)))